# Perpendicular ion heating in turbulence and reconnection: magnetic moment breaking by coherent fluctuations

### v1

@@ 4.1 Balanced turbulence
We will first examine the case of "balanced" turbulence, where the fluxes of Alfvénic fluctuations propagating parallel and antiparallel to the magnetic field are comparable: the imbalanced case is rather different, and is discussed in Sec. 4.3 later. Typical scalings for the fluctuation amplitudes $\delta b_{k}$ and $c\delta E_{k}/B_{0}$ in balanced turbulence are plotted in Figure 2. At small scales, $k_{\perp}\rho_{p}\gg 1$, we have $\delta b_{k}\sim k_{\perp}^{-2/3}$ or steeper (Schekochihin et al., 2009; Boldyrev & Perez, 2012; Zhou et al., 2023); say $\delta b_{k}\sim\delta b_{*}(k_{\perp}\rho_{p})^{-a}$, where $\delta b_{*}$ is then the amplitude at $k_{\perp}\rho_{p}=1$. Then,
@@
At large scales, $k_{\perp}\rho_{p}\ll 1$, the magnetic fluctuations scale as a power law, between $\delta b_{k}\propto k_{\perp}^{-1/3}$ (Goldreich & Sridhar, 1995; Chen et al., 2011) and $\delta b_{k}\propto k_{\perp}^{-1/4}$ (Boldyrev, 2006; Mallet et al., 2016; Chen et al., 2020), where the exact scaling likely depends on the regime of turbulence. Then, $Q_{\perp p}$ is an increasing function of $k_{\perp}$. In balanced turbulence, there is typically a smooth join between the $k_{\perp}\rho_{p}\ll 1$ and $k_{\perp}\rho_{p}\gg 1$ spectra, and so we expect the heating rate as a function of $k_{\perp}$ to reach a maximum at relatively small scales, when $\omega\sim\Omega_{p}$. This agrees (qualitatively at least) with the peak of the heating rate observed in the hybrid-kinetic numerical simulations of Arzamasskiy et al. (2019). We have plotted the proton heating rate (56) as a function of (inverse) scale $k$ in Figure 3, for several different outer scale amplitudes $A=\delta b_{L}/v_{\mathrm{A}}$, assuming that $L/\rho_{p}=10^{4}$ and $v_{th\rm p}/v_{\mathrm{A}}=0.1$. The energy flux into the turbulent cascade is defined as $\epsilon=\delta b_{L}^{3}/L$. The trend agrees with our analysis above: if the amplitude is high enough that $Q_{\perp p}/\epsilon\sim 1$ at $k_{\perp}\rho_{p}\sim 1$, the peak heating is at the proton gyroradius scale (e.g. the red line in Fig. 3). At lower amplitudes, the heating is less efficient, and peaks at a smaller scale (e.g. the blue line in Fig. 3). The oscillations in $Q_{\perp p}$ when $k_{\perp}\rho_{p}>1$ are due to the Bessel function; in reality, structures will have contributions from a range of $k_{\perp}\sim 1/\lambda$ and this behaviour will be smoothed out.
@@ 4.2 Proton heating fraction in balanced turbulence
where $v_{\mathrm{A}}$ is the Alfvén velocity. Assuming $\delta b_{k}\propto k_{\perp}^{-b}$ for $k_{\perp}\rho_{i}\ll 1$ (with $b=1/3$ corresponding to Goldreich & Sridhar (1995) and $b=1/4$ corresponding to the Boldyrev (2006) including dynamic alignment), we have
@@ 4.3 Forced imbalanced Alfvénic turbulence and the helicity barrier
Meyrand et al. (2021) found that in numerical simulations of forced imbalanced turbulence, the helicity barrier causes energy to build up at scales larger than $\rho_{p}$ with a spectral break to a very steep ($\delta b_{k}\propto k_{\perp}^{-3/2}$ or steeper) spectrum beyond $k_{h}\lesssim 1/\rho_{p}$ in the "transition range", before returning to a shallower spectrum at around $k_{\perp}\rho_{p}\sim 1$. We can study this situation with our heating rate expression (57) also. Writing $\delta b_{k}\sim\delta b_{h}(k_{\perp}/k_{h})^{-3/2}$, and taking $k_{\perp}\rho_{p}\lesssim 1$, the frequency $\omega\sim k_{\perp}\delta b_{k}\propto k_{\perp}^{-1/2}$ in the transition range, a decreasing function of $k$. Inserting into our expression for $Q_{\perp p}$ for $k_{\perp}\rho_{p}\ll 1$, we find
@@
Physical systems of imbalanced turbulence, for example the solar wind, are not in fact typically forced, and so the helicity barrier may not cause as sharp a transition to ion heating as suggested on the basis of forced numerical simulations (Meyrand et al., 2021; Squire et al., 2021). The scaling for the maximum heating rate (71) still applies, since the helicity barrier does still cause a steep transition range at scales larger than $\rho_{p}$.
@@ 4.4 Intermittency
So far, we have assumed that the turbulence is characterized by a single amplitude at each scale, $\delta b_{k}$. In reality, $\delta b_{k}$ is a random variable at each scale, typically with a heavy large-amplitude tail; both in solar wind observations (Salem et al., 2009; Zhdankin et al., 2012; Sioulas et al., 2022) and in numerical simulations (Mallet et al., 2015, 2016; Zhdankin et al., 2016a, b). This can dramatically increase the ion heating fraction.
@@
It is worth mentioning that intermittency models including dynamic alignment (e.g. Chandran et al. (2015); Mallet & Schekochihin (2017)) do not necessarily increase the ion heating fraction, because the dynamic alignment between $\delta\boldsymbol{z}^{\pm}$ increases the nonlinear timescale of the turbulent structures, decreasing the effectiveness of the heating due to the exponential suppression factor. This behavior is the opposite to what was found earlier in Mallet et al. (2019), who used an exponential suppression factor based solely on the amplitude (Chandran et al., 2010), rather than the rate of change of the fluctuations.
A more promising intermittency model in this regard (which nevertheless obtains the same 3/2-3/2 perpendicular spectrum) is the reflection-driven turbulence model in Chandran et al. (2025), in which dynamic alignment does not play a role. There, larger-amplitude fluctuations naturally have higher frequencies, as required to make the ion heating more effective. Another approach, independent of any particular intermittency model, would be to use in situ measurements of the distribution of turbulent fluctuations to calculate the heating rate $\langle Q_{\perp}\rangle$ directly. Parker Solar Probe has recently started to explore the plasma environment very close to the sun (Kasper et al., 2021) where ion heating is thought to be particularly important (Kasper & Klein, 2019), and it will be interesting to assess the ion heating in this newly-explored regime.
@@ 4.5 Minor ions
Since $Z_{i}m_{p}/m_{i}<1$, the exponential suppression is less effective for minor ions, exponentially increasing the heating rate relative to the protons. This is similar to what happens in the stochastic heating model applied to minor ions (Chandran, 2010). Note that $\rho_{th}=\rho_{p}(v_{thi}/v_{thp})(Z_{i}m_{p}/m_{i})$, so that the gyroradius of the ions is typically larger than that of the protons: however, in the case of imbalanced turbulence where the cascade is already cut off due to the helicity barrier (Sec. 4.3), this does not have a major effect.
@@ 5 Reconnection
Drake et al. (2009a) developed a theory of perpendicular ion heating in reconnection, and in Drake et al. (2009b) showed that for guide-field reconnection, there was a threshold for strong ion heating by different ion species, namely
@@
where $R\sim 0.1$ is the normalized reconnection rate $u_{in}/u_{out}$, $v_{Ay}$ is the Alfvén speed based on the reconnecting field $\delta B_{y}$, and $v_{\mathrm{A}}$ is the Alfvén speed based on the guide-field $B_{0}$. This was derived in the following way. Given a current sheet width of order $\rho_{s}=c_{s}/\Omega_{p}$, an inflow speed $u_{in}\sim Rv_{Ay}$, the transit time of an ion from the inflow out of the sheet is of order $\tau\sim\rho_{s}/u_{in}$. Comparing this with the ion’s gyrofrequency $\Omega_{i}=\Omega_{p}Z_{i}m_{p}/m_{i}$ and requiring $\tau\Omega_{i}<1$ gives (77). Below the threshold, the ions conserve the first adiabatic invariant $\mu$, while above the threshold, $\mu$ is not conserved and strong ion heating occurs. There is an obvious equivalence between the threshold in $\eta_{K}$ in our estimated heating rate (45) and the Drake et al. (2009b) theory. To make this more concrete, we note that in the current sheet, there is an $E_{x}$ driving the Alfvénic exhaust,
@@
For $\eta\sim\eta_{K}\sim\text{const.}$ (we assume this does not vary with $K$ since the outflow is coherent in time), we use the same argument as Drake et al. (2009a), leading to
@@
Apart from numerical prefactors of order unity which we have not calculated, this agrees with Eq. 6 of Drake et al. (2009a), both in the threshold $\eta\sim 1$ and in the order-of-magnitude of the saturated total heating when the threshold is attained. It would also technically be possible to calculate the heating at a reconnection event using our model by specifying the functional form of the electromagnetic fields precisely, arriving at a more accurate estimate; we will leave this for future work.
This section amounts in some ways to a rederivation of the Drake et al. (2009b) model. However, using our approach, it is perhaps more obvious why the acceleration of the ion into the outflow $E\times B$ velocity is accompanied by ion heating. Drake et al. (2009a) suggest that "the reflected particles interpenetrate with particles that have already crossed the boundary of the exhaust but have not passed through the reversal region", thus gaining an effective temperature. In our approach it is obvious that the diffusion is due to the random initial phase of the particle as it enters the reversal region of the exhaust. While the energy change of each ion $\Delta$ is, on average, zero, considering the plasma distribution function as a whole, the ions acquire a broad range of energies.
@@ 6 Discussion
We have analysed the interaction of an ion with a localized, coherent fluctuation, deriving the change in the ion’s perpendicular and parallel kinetic energy. To lowest order, the energy change is described by Eq. (1), which depends linearly on the amplitude of the fluctuating fields and on a factor $\exp(-1/\eta)$, where $\eta\sim 1/\tau\Omega_{i}$, with $\tau$ the characteristic timescale of the fluctuation and $\Omega_{i}$ the ion’s gyrofrequency. For the whole population of ions, the interaction leads to diffusion in energy and heating. This leads to weak heating for $\eta\ll 1$, and strong heating for $\eta\sim 1$: this reflects the well-known conservation of the magnetic moment for $\eta\ll 1$, and is similar to previous theories of stochastic heating (Chandran et al., 2010). As part of our derivation, we have recovered the fact that, if the fluctuation is a wavepacket with a fixed phase velocity $v_{\mathrm{ph}}$, the diffusion is along circular scattering contours in $_{\parallel}-vv_{\perp}$ space, centered on $v_{\mathrm{ph}}$. (Sec. 3.3), a result usually associated with quasilinear cyclotron resonant heating (Kennel & Engelmann, 1966). Thus, our results combine the physics of stochastic heating and cyclotron heating in a single theoretical framework, based on the interaction of ions with coherent structures. In many systems, our approach based on individual localized fluctuations may be more appropriate than the assumption of infinite plane waves usually required to derive the results of quasilinear theory, for example in strongly nonlinear turbulence and reconnection. The model is also quite easy to apply to different physical situations: one needs estimates for a characteristic timescale relative to the gyroperiod, $1/\eta$, as well as an estimate of the fluctuation amplitude $\varepsilon$ of the structure.
@@
We have applied our results to low-$\beta$ Alfvénic turbulence, obtaining a simple expression for the fraction of the turbulent flux absorbed by the protons in both balanced turbulence and in imbalanced turbulence with a helicity barrier (Meyrand et al., 2021), as may be present in the fast Alfvénic solar wind (Bowen et al., 2020; McIntyre et al., 2024). Our theoretical estimate for the proton heating fraction compares well with the numerically observed estimate for the ion heating fraction in Matthaeus et al. (2016). We also show that our model is well-suited to incorporating different models of intermittency, which we predict should enhance the ion heating: one could also use the observational data to assess this enhancement directly. Moreover, we have determined how the perpendicular lengthscale of the fluctuation affects the heating rates: at smaller amplitudes, the heating does not peak at $k_{\perp}\rho_{\rm th}\sim 1$, but at a smaller scale, as was previously observed in the hybrid-kinetic simulations of Arzamasskiy et al. (2019). We also show that the heating of minor ions is greatly enhanced over the proton heating, another property associated with both cyclotron (Kasper et al., 2013) and stochastic (Chandran, 2010) heating. Our framework can also model ion heating in reconnection, naturally producing perpendicular heating with a threshold similar to the theory and numerical simulations of Drake et al. (2009b, a). The similarity of our results for heating due to reconnection and turbulence suggest that the disparity between these two paradigms for coronal heating (Klimchuk, 2015; Chandran & Hollweg, 2009) may not be as drastic as often thought.
@@
Johnston et al. (2025) have shown within the framework of quasilinear theory that it is possible to recover an exponential or exponential-like suppression of heating, similar to our model and to the original Chandran et al. (2010) stochastic heating theory. Moreover, they show that test-particle heating in both balanced and imbalanced turbulence simulations seems to be well described by such an exponential suppression factor. In both Johnston et al. (2025) and this paper, the exponential factor depends on the typical frequency or inverse timescale of the interactions compared to the ion cyclotron frequency, i.e. $\eta\sim 1/\tau\Omega_{i}$, reinforcing the fact that the relevant physics in both cases is the breaking of the magnetic moment conservation. However, the source of the exponential suppression factor is different: in Johnston et al. (2025), it is due to the exponentially small fraction of turbulent fluctuations at high frequency, whereas in our case, it comes from the fact that $\mu$ is an adiabatic invariant as $\eta\to 0$, conserved to all orders.
@@
Acknowledgements.– AM was supported by NASA grants 80NSSC21K0462 and 80NSSC21K1766. BC was supported in part by NASA under grant number 80NSSC24K0171. KGK was partially supported by NASA grant 80NSSC24K0171 and contract 80ARC021C0001. TE acknowledges funding from The Chuck Lorre Family Foundation Big Bang Theory Graduate Fellowship and NASA grant 80NSSC20K1285. All authors also acknowledge support from NASA contract NNN06AA01C. AM thanks J. Squire, Z. Johnston, and M. Kunz for many useful conversations.

### v2

@@ 4.1 Balanced turbulence
We will first examine the case of "balanced" turbulence, where the fluxes of Alfvénic fluctuations propagating parallel and antiparallel to the magnetic field are comparable: the imbalanced case is rather different, and is discussed in Sec. 4.3 later. Typical scalings for the fluctuation amplitudes $\delta b_{k}$ and $c\delta E_{k}/B_{0}$ in balanced turbulence are plotted in Figure 2. At small scales, $k_{\perp}\rho_{p}\gg 1$, we have $\delta b_{k}\sim k_{\perp}^{-2/3}$ or steeper (Schekochihin et al., 2009; Boldyrev and Perez, 2012; Zhou et al., 2023); say $\delta b_{k}\sim\delta b_{*}(k_{\perp}\rho_{p})^{-a}$, where $\delta b_{*}$ is then the amplitude at $k_{\perp}\rho_{p}=1$. Then,
@@
At large scales, $k_{\perp}\rho_{p}\ll 1$, the magnetic fluctuations scale as a power law, between $\delta b_{k}\propto k_{\perp}^{-1/3}$ (Goldreich and Sridhar, 1995; Chen et al., 2011) and $\delta b_{k}\propto k_{\perp}^{-1/4}$ (Boldyrev, 2006; Mallet et al., 2016; Chen et al., 2020), where the exact scaling likely depends on the regime of turbulence. Then, $Q_{\perp p}$ is an increasing function of $k_{\perp}$. In balanced turbulence, there is typically a smooth join between the $k_{\perp}\rho_{p}\ll 1$ and $k_{\perp}\rho_{p}\gg 1$ spectra, and so we expect the heating rate as a function of $k_{\perp}$ to reach a maximum at relatively small scales, when $\omega\sim\Omega_{p}$. This agrees (qualitatively at least) with the peak of the heating rate observed in the hybrid-kinetic numerical simulations of Arzamasskiy et al. (2019) and Cerri et al. (2021). We have plotted the proton heating rate (56) as a function of (inverse) scale $k$ in Figure 3, for several different outer scale amplitudes $A=\delta b_{L}/v_{\mathrm{A}}$, assuming that $L/\rho_{p}=10^{4}$ and $v_{th\rm p}/v_{\mathrm{A}}=0.1$. The energy flux into the turbulent cascade is defined as $\epsilon=\delta b_{L}^{3}/L$. The trend agrees with our analysis above: if the amplitude is high enough that $Q_{\perp p}/\epsilon\sim 1$ at $k_{\perp}\rho_{p}\sim 1$, the peak heating is at the proton gyroradius scale (e.g. the red line in Fig. 3). At lower amplitudes, the heating is less efficient, and peaks at a smaller scale (e.g. the blue line in Fig. 3). The oscillations in $Q_{\perp p}$ when $k_{\perp}\rho_{p}>1$ are due to the Bessel function; in reality, structures will have contributions from a range of $k_{\perp}\sim 1/\lambda$ and this behaviour will be smoothed out.
@@ 4.2 Proton heating fraction in balanced turbulence
where $v_{\mathrm{A}}$ is the Alfvén velocity. Assuming $\delta b_{k}\propto k_{\perp}^{-b}$ for $k_{\perp}\rho_{i}\ll 1$ (with $b=1/3$ corresponding to Goldreich and Sridhar (1995) and $b=1/4$ corresponding to the Boldyrev (2006) including dynamic alignment), we have
@@ 4.3 Forced imbalanced Alfvénic turbulence and the helicity barrier
Meyrand et al. (2021) found that in numerical simulations of forced imbalanced turbulence, the helicity barrier causes energy to build up at scales larger than $\rho_{p}$ with a spectral break to a very steep ($\delta b_{k}\propto k_{\perp}^{-3/2}$ or steeper) spectrum beyond $k_{h}\sim 2(1-\sigma_{c})^{1/4}/\rho_{p}\lesssim 1/\rho_{p}$ in the "transition range", before returning to a shallower spectrum at around $k_{\perp}\rho_{p}\sim 1$. We can study this situation with our heating rate expression (57) also. Writing $\delta b_{k}\sim\delta b_{h}(k_{\perp}/k_{h})^{-3/2}$, and taking $k_{\perp}\rho_{p}\lesssim 1$, the frequency $\omega\sim k_{\perp}\delta b_{k}\propto k_{\perp}^{-1/2}$ in the transition range, a decreasing function of $k$. Inserting into our expression for $Q_{\perp p}$ for $k_{\perp}\rho_{p}\ll 1$, we find
@@
Physical systems of imbalanced turbulence, for example the solar wind, are not in fact homogeneous nor forced in the usual way, and so the helicity barrier may not have sufficient time to reach steady state and cause as sharp a transition to ion heating as suggested on the basis of forced numerical simulations (Meyrand et al., 2021; Squire et al., 2021). The scaling for the maximum heating rate (71) still applies, since the helicity barrier does still cause a steep transition range at scales larger than $\rho_{p}$.
@@ 4.4 Intermittency
So far, we have assumed that the turbulence is characterized by a single amplitude at each scale, $\delta b_{k}$. In reality, $\delta b_{k}$ is a random variable at each scale, typically with a heavy large-amplitude tail; both in solar wind observations (Salem et al., 2009; Zhdankin et al., 2012; Sioulas et al., 2022) and in numerical simulations (Mallet et al., 2015, 2016; Zhdankin et al., 2016a, b). This can dramatically increase the ion heating fraction (Mallet et al., 2019; Cerri et al., 2021).
@@
It is worth mentioning that intermittency models including dynamic alignment (e.g. Chandran et al. (2015); Mallet and Schekochihin (2017)) do not necessarily increase the ion heating fraction, because the dynamic alignment between $\delta\boldsymbol{z}^{\pm}$ increases the nonlinear timescale of the turbulent structures, decreasing the effectiveness of the heating due to the exponential suppression factor. This decrease in ion heating fraction due to dynamic alignment is the opposite to what was found earlier in Mallet et al. (2019), who used an exponential suppression factor based solely on the amplitude (Chandran et al., 2010), rather than the rate of change of the fluctuations, thus finding that dynamic alignment acted to increase the stochastic heating rate.
A more promising intermittency model in this regard (which nevertheless obtains the same 3/2-3/2 perpendicular spectrum) is the reflection-driven turbulence model in Chandran et al. (2025), in which dynamic alignment does not play a role. There, larger-amplitude fluctuations naturally have higher frequencies, as required to make the ion heating more effective. Another approach, independent of any particular intermittency model, would be to use in situ measurements of the distribution of turbulent fluctuations to calculate the heating rate $\langle Q_{\perp}\rangle$ directly. Parker Solar Probe has recently started to explore the plasma environment very close to the sun (Kasper et al., 2021) where ion heating is thought to be particularly important (Kasper and Klein, 2019), and it will be interesting to assess the ion heating in this newly-explored regime.
@@ 4.5 Minor ions
Since $Z_{i}m_{p}/m_{i}<1$, the exponential suppression is less effective for minor ions, exponentially increasing the heating rate relative to the protons. This is similar to what happens in the stochastic heating model applied to minor ions (Chandran, 2010). Note that $\rho_{th}=\rho_{p}(v_{thi}/v_{thp})(Z_{i}m_{p}/m_{i})$, so that the gyroradius of the ions is typically larger than that of the protons: however, in the case of imbalanced turbulence where the cascade is already cut off due to the helicity barrier (Sec. 4.3), this may have less of an effect. An exception is in forced homogeneous numerical simulations, where the saturated state of the turbulence has such a high amplitude that fluctuations attain the minor ion’s gyrofrequency at large scales, causing extreme heating of minor species (Zhang et al., 2025).
@@ 5 Reconnection
Drake et al. (2009b) developed a theory of perpendicular ion heating in reconnection, and in Drake et al. (2009a) showed that for guide-field reconnection, there was a threshold for strong ion heating by different ion species, namely
@@
where $R\sim 0.1$ is the normalized reconnection rate $u_{in}/u_{out}$, $v_{Ay}$ is the Alfvén speed based on the reconnecting field $\delta B_{y}$, and $v_{\mathrm{A}}$ is the Alfvén speed based on the guide-field $B_{0}$. This was derived in the following way. Given a current sheet width of order $\rho_{s}=c_{s}/\Omega_{p}$, an inflow speed $u_{in}\sim Rv_{Ay}$, the transit time of an ion from the inflow out of the sheet is of order $\tau\sim\rho_{s}/u_{in}$. Comparing this with the ion’s gyrofrequency $\Omega_{i}=\Omega_{p}Z_{i}m_{p}/m_{i}$ and requiring $\tau\Omega_{i}<1$ gives (77). Below the threshold, the ions conserve the first adiabatic invariant $\mu$, while above the threshold, $\mu$ is not conserved and strong ion heating occurs. There is an obvious equivalence between the threshold in $\eta_{K}$ in our estimated heating rate (45) and the Drake et al. (2009a) theory. To make this more concrete, we note that in the current sheet, there is an $E_{x}$ driving the Alfvénic exhaust,
@@
For $\eta\sim\eta_{K}\sim\text{const.}$ (we assume this does not vary with $K$ since the outflow is coherent in time), we use the same argument as Drake et al. (2009b), leading to
@@
Apart from numerical prefactors of order unity which we have not calculated, this agrees with Eq. 6 of Drake et al. (2009b), both in the threshold $\eta\sim 1$ and in the order-of-magnitude of the saturated total heating when the threshold is attained. It would also technically be possible to calculate the heating at a reconnection event using our model by specifying the functional form of the electromagnetic fields precisely, arriving at a more accurate estimate; we will leave this for future work.
This section amounts in some ways to a rederivation of the Drake et al. (2009a) model. However, using our approach, it is perhaps more obvious why the acceleration of the ion into the outflow $E\times B$ velocity is accompanied by ion heating. Drake et al. (2009b) suggest that "the reflected particles interpenetrate with particles that have already crossed the boundary of the exhaust but have not passed through the reversal region", thus gaining an effective temperature. In our approach it is obvious that the diffusion is due to the random initial phase of the particle as it enters the reversal region of the exhaust. While the energy change of each ion $\Delta$ is, on average, zero, considering the plasma distribution function as a whole, the ions acquire a broad range of energies.
@@ 6 Discussion
We have analysed the interaction of an ion with a localized, coherent fluctuation, deriving the change in the ion’s perpendicular and parallel kinetic energy. To lowest order, the energy change is described by Eq. (1), which depends linearly on the amplitude of the fluctuating fields and on a factor $\exp(-1/\eta)$, where $\eta\sim 1/\tau\Omega_{i}$, with $\tau$ the characteristic timescale of the fluctuation and $\Omega_{i}$ the ion’s gyrofrequency. For the whole population of ions, the interaction leads to diffusion in energy and heating. This leads to weak heating for $\eta\ll 1$, and strong heating for $\eta\sim 1$: this reflects the well-known conservation of the magnetic moment for $\eta\ll 1$, and is similar to previous theories of stochastic heating (Chandran et al., 2010). As part of our derivation, we have recovered the fact that, if the fluctuation is a wavepacket with a fixed phase velocity $v_{\mathrm{ph}}$, the diffusion is along circular scattering contours in $_{\parallel}-vv_{\perp}$ space, centered on $v_{\mathrm{ph}}$. (Sec. 3.3), a result usually associated with quasilinear cyclotron resonant heating (Kennel and Engelmann, 1966). Thus, our results combine the physics of stochastic heating and cyclotron heating in a single theoretical framework, based on the interaction of ions with coherent structures. In many systems, our approach based on individual localized fluctuations may be more appropriate than the assumption of infinite plane waves usually required to derive the results of quasilinear theory, for example in strongly nonlinear turbulence and reconnection. The model is also quite easy to apply to different physical situations: one needs estimates for a characteristic timescale relative to the gyroperiod, $1/\eta$, as well as an estimate of the fluctuation amplitude $\varepsilon$ of the structure.
@@
We have applied our results to low-$\beta$ Alfvénic turbulence, obtaining a simple expression for the fraction of the turbulent flux absorbed by the protons in both balanced turbulence and in imbalanced turbulence with a helicity barrier (Meyrand et al., 2021), as may be present in the fast Alfvénic solar wind (Bowen et al., 2020; McIntyre et al., 2024). Our theoretical estimate for the proton heating fraction compares well with the numerically observed estimate for the ion heating fraction in Matthaeus et al. (2016). We also show that our model is well-suited to incorporating different models of intermittency, which we predict should enhance the ion heating: one could also use the observational data to assess this enhancement directly. Moreover, we have determined how the perpendicular lengthscale of the fluctuation affects the heating rates: at smaller amplitudes, the heating does not peak at $k_{\perp}\rho_{\rm th}\sim 1$, but at a smaller scale, as was previously observed in the hybrid-kinetic simulations of Arzamasskiy et al. (2019). We also show that the heating of minor ions is greatly enhanced over the proton heating, another property associated with both cyclotron (Kasper et al., 2013) and stochastic (Chandran, 2010) heating. Our framework can also model ion heating in reconnection, naturally producing perpendicular heating with a threshold similar to the theory and numerical simulations of Drake et al. (2009a, b). The similarity of our results for heating due to reconnection and turbulence suggest that the disparity between these two paradigms for coronal heating (Klimchuk, 2015; Chandran and Hollweg, 2009) may not be as drastic as often thought.
@@
Johnston et al. (2025) have shown within the framework of quasilinear theory that it is possible to recover an exponential or exponential-like suppression of heating, similar to our model and to the original Chandran et al. (2010) stochastic heating theory. Moreover, they show that test-particle heating in both balanced and imbalanced turbulence simulations seems to be well described by such an exponential suppression factor. In both Johnston et al. (2025) and this paper, the exponential factor depends on the typical frequency or inverse timescale of the interactions compared to the ion cyclotron frequency, i.e. $\eta\sim 1/\tau\Omega_{i}$, reinforcing the fact that the relevant physics in both cases is the breaking of the magnetic moment conservation. However, the source of the suppression factor is different: in Johnston et al. (2025), the suppression factor in imbalanced turbulence is due to the steep $k_{\parallel}$ spectrum outside the critical balance cone leading to very little power at the high resonant frequency, whereas in our case, it comes from the fact that $\mu$ is an adiabatic invariant as $\eta\to 0$, conserved to all orders.
@@
Acknowledgements.– AM was supported by NASA grants 80NSSC21K0462 and 80NSSC21K1766. BC was supported in part by NASA under grant number 80NSSC24K0171. KGK was partially supported by NASA grant 80NSSC24K0171 and contract 80ARC021C0001. TE acknowledges funding from The Chuck Lorre Family Foundation Big Bang Theory Graduate Fellowship and NASA grant 80NSSC20K1285. All authors also acknowledge support from NASA contract NNN06AA01C. AM thanks J. Squire, Z. Johnston, and M. Kunz for many useful conversations, and the three referees for insightful comments, which significantly improved this paper.